OpenStudy (anonymous):

Which represents the type of sequence: -99, -96, -92, -87, -81, …? (Points : 3) both arithmetic and geometric neither arithmetic nor geometric geometric arithmetic

OpenStudy (the_fizicx99):

Um, I don't think it's neither arithmetic or geometric. -96/-99 = 0.969696... -92/-96 = 0.958... -81 / -81 = 0.9310... --------------------- -96 - (-99) = -96 + 99 = 3 -92 - (-96) = -92 + 96 = 4 -87 - (-92) = -87 + 92 = 5

OpenStudy (the_fizicx99):

Typo, -81 / -87 = 0.9310...

OpenStudy (anonymous):

So then what would it be?

OpenStudy (the_fizicx99):

tHe_FiZiCx99: Um, I don't think it's neither arithmetic or geometric.

OpenStudy (the_fizicx99):

What I meant was, it's \(\ not \) arithmetic or geometric.

OpenStudy (the_fizicx99):

So b >.>